OpenStudy (anonymous):

satellite please explain how you got your answer from last post The formula S = 16t^2 + v0t + s0 tells the height at time t of a projectile fired vertically from height s0 with initial velocity v0. Suppose a projectile is fired vertically from ground level at an initial speed of 448 ft per second. When will the projectile be higher than 3072 ft?

OpenStudy (amistre64):

-16t^2 ....

OpenStudy (amistre64):

mathematics doesnt tell you how to find out physics; it just helps you to express what you find

OpenStudy (amistre64):

the acceleration due to gravity is a constant; it stays the same. on earth that constant is -32 ft/sec/sec

OpenStudy (anonymous):

solve for the equation\[3072\leq-16t^2+v_0t +s_0\ |_{v_0 = 448,\ \ \ s_0=0}\]

OpenStudy (amistre64):

a(t) = -32 it just so happens that \(\int a(t)\) = velocity v(t) = -32t + C

OpenStudy (anonymous):

oops I mean 3072 < expression

OpenStudy (anonymous):

ok lets start at the beginning without integration etc

OpenStudy (amistre64):

lets call it better as: v(t) = -32t + Vi ; for initial velocity at time = 0 another neat trick is that velocity is the derivative of height/position suc that: \(\int v(t)\) = s(t)+C

OpenStudy (anonymous):

ok lets start at the beginning without integration etc

OpenStudy (amistre64):

or rather: s(t) = -16 t^2 +Vi t + C; where the initial position at time = 0 is given as C such that: s(t) = -16 t^2 +Vi t + Hi

OpenStudy (anonymous):

good god is this slow

OpenStudy (amistre64):

its been like that all day, i emailed the shadow and he might be on in a few hours to see what he can do about it

OpenStudy (anonymous):

print filter(None,map(lambda y:y*reduce(lambda x,y:x*y!=0, map(lambda x,y=y:y%x,range(2,int(pow(y,0.5)+1))),1),range(2,1000)))

OpenStudy (anonymous):

ok you have the formula given to you above, and we see that the initial height was 0 because it was launched from the ground and the initial velocity was 448 because you are told

OpenStudy (anonymous):

oops I mean I got 12s < t < 16s

OpenStudy (anonymous):

so you get the formula \[S(t)=-16t^2+448t\] and you are asked when this is greater than 3072, so put \[-16t^2+448t>3072\] and solve

OpenStudy (anonymous):

ok the problem is i dont know how to solve it :/

OpenStudy (anonymous):

To solve it, use the quadratic formula, or complete the square

OpenStudy (anonymous):

first step is \[-16t^2+448t-3702>0\] second step is \[-16(t-12)(t-16)>0\]

OpenStudy (anonymous):

damn is this slow. ok almost done. the zeros are at t=12 and t=16, and it is positive between them, so that is your answer

OpenStudy (anonymous):

positive between them because this is a parabola that faces down, so it is positive between the zeros and negative outside. there finished!

OpenStudy (anonymous):

sorry it took so damned long but for some reason i cannot load the pages

OpenStudy (anonymous):

thank you all so much i guess i just dont unerstand that whole parabola thing and the whole 12<x<16 or whatever it was